OpenStudy (anonymous):

log (base 6) (x+3) + log (base 6) (x+4)=1 **open question to see the equation properly.

OpenStudy (anonymous):

\[\log_{6}(x+3)+\log_{6}(x+4)=1 \]

OpenStudy (anonymous):

Basically, let both sides be exponents of \(6\).

OpenStudy (anonymous):

\[6^{\log_{6}(x+3)+\log_{6}(x+4)}=6^1 \]

OpenStudy (anonymous):

what does that do though?

OpenStudy (anonymous):

It will get rid of the \(\log_6\) if you know your properties of exponents.

OpenStudy (anonymous):

Do you know\[ a^{b+c} =\ldots \]

OpenStudy (anonymous):

would that make it (x+3)(x+4)=6?

OpenStudy (anonymous):

... Yes

OpenStudy (anonymous):

alright one sec let me finish working i tout and see if it works out

OpenStudy (anonymous):

it worked thanks!

OpenStudy (anonymous):

\[9^x+4 \times 3^x-3=0\] @wio could you help me with this problem too please?

OpenStudy (anonymous):

\[9^x+4 \times 3^x-3=0\] Well, let \(u=3^x\).

OpenStudy (anonymous):

\[ u^2+4u-3=0 \]

OpenStudy (anonymous):

Our \(x\) solutions correspond to our \(u\) solutions.

OpenStudy (anonymous):

I mean, we could have alternatively substituted \(x =\log_3(u)\).

OpenStudy (anonymous):

the answer given is in log form so does that mean i use the alternative?

OpenStudy (jhannybean):

\[\log(a+b)=\log(ab)\]\[\log_6(x+3) +\log_6(x+4) =1\]\[\log_6[(x+3)(x+4)] =1\]\[(x+3)(x+4)=6\]

OpenStudy (jhannybean):

That's how I tried solving it.

OpenStudy (anonymous):

that's kind of what i got

OpenStudy (anonymous):

\[\log(a+b)=\log(ab)\]???

OpenStudy (anonymous):

No, what it means is you solve for \(u\). You use quadratic formula.

OpenStudy (anonymous):

Then, ones you have your two values for \(u\)... suppose they are \(u_1,u_2\), then you solve for \(x\)\[ 3^{x_1}=u_1\\ 3^{x_2}=u_2\\ \]

OpenStudy (anonymous):

In short, find your \(u\) values, and find your \(x\) values using: \(x = \log_3(u)\).

OpenStudy (anonymous):

\[ u^2+4u-3=0 \]So :\[ u = \frac{-4\pm\sqrt{4^2-4(1)(-3)}}{2(1)} \]And \[ x = \log_3\left(\frac{-4\pm\sqrt{4^2-4(1)(-3)}}{2(1)}\right) \]

OpenStudy (anonymous):

\wait when you solve that what ar eyou supposed to get at the end?

OpenStudy (anonymous):

@wio